# Diverse, not Short: A Length-Controlled Self-Learning Framework for Improving Response Diversity of Language Models

### v2

@@ 6 Results
In our CWT evaluations (Tab. 3), Diverse-NS produces the highest DSI scores for both Llama-8B and Olmo-7B. Interestingly, for Llama-8B the other methods actually reduce the ArmoRM score below baseline but Diverse-NS exceeds it. The highest 4-gram diversity is observed for Diverse-NS or -Lite in all cases. We also compute with six lexical diversity measures and ArmoRM. Both Diverse-NS and its lightweight variant significantly outperform DivPO on every diversity metric. The remains above the baseline for all metrics except MAAS, where it dips marginally below and similarly shows a slight under-performance for ArmoRM. Crucially, even where suggests a minor quality drop, the absolute diversity values after self-training still exceed those of the base model (despite longer outputs), indicating that any loss in writing quality is minimal (refer to Appendix F for Llama-8B and Olmo-13B results)555We provide all results with std. dev. values in Table E.1.
@@ 7 Discussion
We introduced Diverse-NS, a self-learning framework to improve output diversity while preserving quality. Experiments with Llama-8B and Olmo-7B show that Diverse-NS improves diversity on four creative generation tasks: DAT, PGT, AUT, CWT.
@@
Overall, Diverse-NS offers a practical and scalable solution for boosting diversity in aligned LLMs. By addressing the length bias in both training and evaluation, our framework sets a foundation for more expressive and diverse language generation. We hope this work encourages further exploration of length-aware diversity alignment.
@@
While our study demonstrates the effectiveness of diversity-aware self-learning, several areas remain open for future exploration. First, our data filtering relies on a single diversity metric (e.g., entropy or TTR). Although effective, no single metric can fully capture all aspects of text diversity. Future work could incorporate multiple metrics to jointly optimize lexical, semantic, and syntactic variation, as well as novelty, to better capture diverse training signals. Second, we focus on one data generation task—short story writing—which allows for controlled analysis and task-specific improvements. Expanding the framework to include a broader set of tasks could lead to more generalizable diversity enhancements. Third, our self-learning setup investigates only a single round of preference tuning. While this provides a strong baseline, recent work suggests that repeated rounds of self-training can affect diversity (Guo et al., 2023; Seddik et al., 2024; Herel and Mikolov, 2024). It would be valuable to study how diversity evolves across multiple self-learning iterations in our framework. We do not include human evaluation in this study. While human judgments can provide nuanced insight, they often come with variability and inconsistency. Along these lines, it is often prohibitively costly to gather high-quality human feedback—particularly at the scale necessary to provide stable estimates. In this paper, we emphasize stringent empirical evaluation of D-NS using reliable, automatic metrics and leave human-centered evaluation for future work. It is worth noting a peculiar change in the length distribution of the preference-tuning model (Table D.1). Even though preference pairs are of comparable lengths in Diverse-NS and Diverse-NS-Lite, the model learns to be more expressive. We suspect this shift is influenced by a skewed proportion of longer preference pairs, which may inadvertently bias the model toward generating longer responses. Controlling the length distribution is challenging under our current framework due to the strict filtering criteria. In future work, we aim to address this by extending our method to a multi-task setup that includes both short and long generation tasks.
@@ Appendix B Pilot Analysis for Sequential Prompting
We conducted an exploratory analysis on short stories generated from Llama-3.1-8B and Olmo-2-7B models Grattafiori et al. (2024); OLMo et al. (2024). The analysis was targeted at understanding the repeating patterns in the generated stories. With the help of the diversity package in Python Shaib et al. (2024a), we extract the top-5 repeating Part-Of-Speech (POS) bi-grams. We find that the most repeated bigram (IN DT) occurs in over 15k stories (out of 20k) and of occurrences are present at the beginning of the generated story, refer to table B.1.
@@
Based on the findings, we conducted a sequential prompting experiment that elicits a more diverse response from the model by asking the model to avoid repeating phrases (refer to section A.1 for exact prompts). We find that the diversity of the second response is, on average, higher than the first one.
@@
Appendix C Hyperparameters for Preference Optimization
@@
Appendix D Reponse Length Distribution
We observe that the length distribution varies after fine-tuning the model. As presented in table D.1, we observe that the average (and standard deviation) of response length reduces for DivPO and increases for our proposed methods (Diverse-NS and Diverse-NS-Lite). DivPO (inadvertently) teaches the model to generate shorter responses (refer to table 2). Despite maintaining comparable length for “chosen” and “rejected” samples in our methods (Diverse-NS and Diverse-NS-Lite), the model interestingly learns to generate longer responses. We suspect this shift is influenced by a skewed proportion of longer preference pairs, which may inadvertently bias the model toward generating longer responses.
@@ Appendix D Reponse Length Distribution
Appendix E Results with Standard Deviation
In this section, we report the results with the standard deviation values in Table E.1.
@@ Appendix E Results with Standard Deviation
Appendix F DD-based Evaluation
Similar to the results presented fig. 1 for Olmo-7B, we present the results for Llama-8B and Olmo-13B in this section.
Appendix G A Summary of Metrics
We provide a concise summary of all metrics used in our evaluation setup in Table G.1.

### v3

@@ 6 Results
In our CWT evaluations (Tab. 3), Diverse-NS produces the highest DSI scores for both Llama-8B and Olmo-7B. Interestingly, for Llama-8B the other methods actually reduce the ArmoRM score below baseline but Diverse-NS exceeds it. The highest 4-gram diversity is observed for Diverse-NS or -Lite in all cases. We also compute with six lexical diversity measures and ArmoRM. Both Diverse-NS and its lightweight variant significantly outperform DivPO on every diversity metric. The remains above the baseline for all metrics except MAAS, where it dips marginally below and similarly shows a slight under-performance for ArmoRM. Crucially, even where suggests a minor quality drop, the absolute diversity values after self-training still exceed those of the base model (despite longer outputs), indicating that any loss in writing quality is minimal (refer to Appendix˜I for Llama-8B and Olmo-13B results)888We provide all results with std. dev. values in Table H.1.
Significance of DSI Improvements.
In line with the findings of Johnson et al. (2023), we observe a narrow distribution of DSI values in our CWT experiments (percentiles: ). Given the narrow distribution, to highlight the significance of the DSI improvements achieved with D-NS or D-NS-Lite, we conduct an independent t-test on method pairs, e.g., base model versus D-NS, with 70 DSI measurements (for 70 unique CWT responses) for each method. Results are presented in Table˜4, where negative t-statistics indicate that “Method B” yields higher response diversity than “Method A”. We find that both D-NS and D-NS-Lite produce significantly more diverse responses than the base model and DivPO. On the other hand, the differences between the base model and DivPO, as well as between D-NS and D-NS-Lite, are not statistically significant. These findings further underscore the effectiveness of D-NS-Lite as a computationally efficient alternative to D-NS.
Method A | Method B | t-statistic | p-value |
LLaMA-8B | |||
Base | DivPO | Not Significant | |
[…]
D-NS | D-NS-Lite | 1.5492 | Not Significant |
Olmo-7B | |||
Base | DivPO | Not Significant | |
[…]
D-NS | D-NS-Lite | Not Significant |
Human Evaluation.
To further understand the human perceived value of DSI improvements, we conduct a human evaluation on the CWT responses. Under pairwise evaluation, annotators are shown a prompt along with two responses from two different methods (e.g., D-NS versus DivPO) and asked to select the more diverse response (see Section˜A.2). Per CWT prompt, we select 20 least similar (Jaccard sim.) pairs to ensure maximally distinguishable comparisons between methods. Using 7 prompts, we construct 140 pairs each for D-NS vs. Base and D-NS vs. DivPO. The resulting 280 comparisons are evenly and randomly assigned to four annotators: two involved in the study (E1, E2) and two independent evaluators with no prior exposure (E3, E4). Human evaluation confirms that D-NS consistently outperforms both baselines in win percentage. Except for E3, all annotators strongly preferred D-NS, with win rates reaching as high as against DivPO (E4). These findings suggest that even modest improvements in DSI scores correspond to clear and consistent gains in human-perceived diversity. We extended our human evaluation to the LLM-as-a-Judge evaluation that covers comparison between D-NS-Lite and the baselines as well. The results of the LLM-as-a-Judge experiment reciprocate the human evaluation findings and underscore the effectiveness of both D-NS and D-NS-Lite (refer to Appendix˜K for details). We provide a few examples of the model responses in Appendix˜L for the readers.
@@ 7 Discussion
We introduced Diverse-NS, a data selection strategy to improve output diversity while preserving quality. Experiments with Llama-8B and Olmo-7B show that Diverse-NS improves diversity on four creative generation tasks: DAT, PGT, AUT, CWT. On CWT, the diversity gains achieved by Diverse-NS and its lightweight variant, Diverse-NS-Lite, are statistically significant. These improvements are further supported by human evaluations, where Diverse-NS achieves higher win rates against both the base model and DivPO.
Human | D-NS vs. Base | D-NS vs. DivPO |
Evaluator | (D-NS Win% - Tie%) | (D-NS Win% - Tie%) |
E1 | ||
E2 | ||
E3 | ||
E4 |
@@
Overall, Diverse-NS offers a practical and scalable solution for boosting diversity in aligned LLMs. By addressing the length bias in both training and evaluation, our sets a foundation for more expressive and diverse language generation. We hope this work encourages further exploration of length-aware diversity alignment.
@@
While our study demonstrates the effectiveness of diversity-aware self-learning, several areas remain open for future exploration. First, our data filtering relies on a single diversity metric (e.g., entropy or TTR). Although effective, no single metric can fully capture all aspects of text diversity. Future work could incorporate multiple metrics to jointly optimize lexical, semantic, and syntactic variation, as well as novelty, to better capture diverse training signals. Second, we focus on one data generation task—short story writing—which allows for controlled analysis and task-specific improvements. Expanding the framework to include a broader set of tasks could lead to more generalizable diversity enhancements. Third, our self-learning setup investigates only a single round of preference tuning. While this provides a strong baseline, recent work suggests that repeated rounds of self-training can affect diversity (Guo et al., 2023; Seddik et al., 2024; Herel and Mikolov, 2024). It would be valuable to study how diversity evolves across multiple self-learning iterations in our framework. It is worth noting a peculiar change in the length distribution of the preference-tuning model (Table˜G.1). Even though preference pairs are of comparable lengths in Diverse-NS and Diverse-NS-Lite, the model learns to be more expressive. We suspect this shift is influenced by a skewed proportion of longer preference pairs, which may inadvertently bias the model toward generating longer responses. Controlling the length distribution is challenging under our current framework due to the strict filtering criteria. In future work, we aim to address this by extending our method to a multi-task setup that includes both short and long generation tasks.
@@
Acknowledgment
This work was funded in part by an Amazon AGI research award to Anna Rumshisky.
@@ A.2 Model Evaluation Prompts
Instruction for Human Evaluators
The instruction provided to the human evaluators is as follows,
Task Description for the LLM-as-a-Judge Evaluation
The task description for the LLM-as-a-Judge evaluation is as follows,
@@ Appendix B Pilot Analysis for Sequential Prompting
We conducted an exploratory analysis on short stories generated from Llama-3.1-8B and Olmo-2-7B models Grattafiori et al. (2024); OLMo et al. (2024). The analysis was targeted at understanding the repeating patterns in the generated stories. With the help of the diversity package in Python Shaib et al. (2024a), we extract the top-5 repeating Part-Of-Speech (POS) bi-grams. We find that the most repeated bigram (IN DT) occurs in over 15k stories (out of 20k) and of occurrences are present at the beginning of the generated story, refer to table˜B.1.
@@
Based on the findings, we conducted a sequential prompting experiment that elicits a more diverse response from the model by asking the model to avoid repeating phrases (refer to section˜A.1 for exact prompts). We find that the diversity of the second response is, on average, higher than the first one.
@@
Appendix C Brief Description of ArmoRM score
ArmoRM uses a Llama-3-8B backbone frozen and attaches 19 “meter” heads that score crowd-rated qualities such as creativity, factuality, safety, and verbosity. A small prompt-conditioned gating network assigns non-negative weights to these meters and sums them, after explicitly removing any correlation with length to avoid verbosity bias. In practice the gate down-weights irrelevant meters: on creative-writing prompts of the weight falls on creativity-related meters, on code-style. The model achieves pairwise accuracy on RewardBench, a benchmark whose preference labels are entirely human-annotated; this already reflects agreement with human judgements.
Appendix D Computational Savings Due to Lightweight Filtration
D-NS-Lite was explicitly designed with practical efficiency in mind. Computing entropy and ArmoRM scores for 400,000 generated sequences incurs two additional forward passes per example—one through the base language model and one through the reward model. With an average response length of 150 tokens, each pass processes approximately 60 million tokens. Assuming inference with a 7B-parameter model999We use Olmo-7B, LLaMA-8B, and ArmoRM, which has approximately 7.5B parameters., this results in an estimated FLOPs per inference round based on the method outlined in Kaplan et al. (2020), totaling around PFLOPs for both passes. This computational cost—spent solely to measure diversity and quality—represents a significant overhead, roughly equivalent to or exceeding the cost of generating the data itself. As model sizes or dataset scales increase, the cost of metric computation will grow rapidly and can quickly become prohibitive. This motivates the need for low-cost alternatives to expensive metrics like entropy and reward model scores. D-NS-Lite addresses this by replacing these metrics with lightweight lexical proxies such as TTR (Type-Token Ratio) and MAAS. These proxies entirely eliminate the need for the two additional inference rounds, resulting in substantial GPU computation savings. Although proxy metrics are computed on the CPU, their cost is negligible and can be handled efficiently even with modest hardware. Notably, in sequential prompting setups (see Section˜4.1), entropy computation for the first set of responses can be avoided, as these are already conditioned on the test-time prompt. However, if one wishes to use later-stage responses to construct the preference dataset, an additional forward pass would be required for each of those responses to match the intended test-time input distribution. By contrast, D-NS-Lite avoids this complication entirely, offering a significantly more efficient pipeline.
Appendix E Implementation of DivPO
Brief Description.
Lanchantin et al. (2025), similar to our study, propose a data selection strategy for preference tuning to enhance response diversity in base models. Their approach involves two main steps for selecting a preference pair from responses to a fixed prompt. First, responses are ranked by a text quality score, with the top forming the pool of “chosen” candidates and the bottom forming the “rejected” pool. Second, each pool is sorted by a diversity metric. A single preference tuning pair is then formed by selecting the most diverse response from the “chosen” pool and the least diverse response from the “rejected” pool.
Hyperparameters.
The key hyperparameters in Lanchantin et al. (2025) are the text quality metric, the diversity metric, and the value of . They experiment with rule-based rewards (e.g., JSON validity) and the ArmoRM reward model as quality metrics. In our study, we use ArmoRM exclusively. While they explore various values, we instead define the “chosen” pool as the top quartile (ArmoRM score percentile) and the “rejected” pool as the bottom quartile (ArmoRM score percentile). For diversity, they use log probabilities, whereas we use entropy.
Appendix F Hyperparameters for Preference Optimization
@@
Appendix G Reponse Length Distribution
We observe that the length distribution varies after fine-tuning the model. As presented in table˜G.1, we observe that the average (and standard deviation) of response length reduces for DivPO and increases for our proposed methods (Diverse-NS and Diverse-NS-Lite). DivPO (inadvertently) teaches the model to generate shorter responses (refer to table˜2). Despite maintaining comparable length for “chosen” and “rejected” samples in our methods (Diverse-NS and Diverse-NS-Lite), the model interestingly learns to generate longer responses. We suspect this shift is influenced by a skewed proportion of longer preference pairs, which may inadvertently bias the model toward generating longer responses.
@@ Appendix G Reponse Length Distribution
Appendix H Results with Standard Deviation
In this section, we report the results with the standard deviation values in Table˜H.1.
@@ Appendix H Results with Standard Deviation
Appendix I DD-based Evaluation
Similar to the results presented fig.˜1 for Olmo-7B, we present the results for Llama-8B and Olmo-13B in this section.
Appendix J A Summary of Metrics
We provide a concise summary of all metrics used in our evaluation setup in Table˜J.1.
@@ Appendix I DD-based Evaluation
Appendix K LLM-as-a-Judge Evaluation on CWT
We extend our human evaluation setup to assess diversity gains using the LLM-as-a-Judge framework. Similar to the human study, we present a single prompt and two model responses to an LLM and ask it to select the more diverse output (see Section˜A.2 for prompts). For each CWT prompt, we generate 10 responses per method, yielding 100 unique pairs for direct comparison. To ensure maximally distinguishable outputs, we select the 50 pairs with the lowest Jaccard similarity. Using 7 prompts, we construct 350 evaluation pairs per method comparison. Each comparison is evaluated by five proprietary LLMs: Claude-3.5-Haiku, Claude-Sonnet-4, GPT-4.1, GPT-4.1-mini, and GPT-4o-mini. We report the win percentages for D-NS and D-NS-Lite against the base model, DivPO, and each other in Table˜K.1. The results align with human evaluation findings: D-NS outperforms both baselines in all but one case, and D-NS-Lite achieves similar improvements, also outperforming the baselines in all but one setting. The comparison between D-NS and D-NS-Lite is more competitive, with win rates hovering around , slightly favoring D-NS.
Judge | D-NS vs. Base | D-NS vs. DivPO | D-NS vs. D-NS-Lite |
LLM | D-NS Win | D-NS Win | D-NS Win |
claude-3-5-haiku | |||
claude-sonnet-4 | |||
gpt-4.1 | |||
gpt-4.1-mini | |||
gpt-4o-mini | |||
Judge | D-NS-Lite vs. Base | D-NS-Lite vs. DivPO | D-NS-Lite vs. D-NS |
LLM | D-NS-Lite Win | D-NS-Lite Win | D-NS-Lite Win |
claude-3-5-haiku | |||
claude-sonnet-4 | |||
gpt-4.1 | |||
gpt-4.1-mini | |||
gpt-4o-mini |
Appendix L Examples of Model Responses for CWT Prompts
In this section, we provide a couple of examples of the model responses to CWT prompts, before and after the diversity tuning. Please refer to the Figure˜L.1.